OpenStudy (anonymous):

For which system of equations would elimination be the simplest technique to choose? A. y = 2x and x + 3y = 5 B. 4x + 3y = 6 and 3y = 5 C. x + y = 7 and x - y = 12 D. 3x - 2y = 8 and 4x - 5y = 20

OpenStudy (anonymous):

B. 4x + 3y = 6 and 3y = 5

OpenStudy (anonymous):

i think it is c not not sure Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

OpenStudy (anonymous):

New thread.

Directrix (directrix):

C) x + y = 7 and x - y = 12 The y variables would add to 0 thus eliminating the y variable. Theory: In the ‘elimination’ method for solving simultaneous equations, two equations are simplified by adding them or subtracting them. This eliminates one of the variables so that the other variable can be found. http://www.teacherschoice.com.au/maths_library/algebra/alg_10.htm